OpenStudy (anonymous):

I need help writing the equalities for this word problem. someone please help?

OpenStudy (anonymous):

A machine is able to cut raw metal sheets in four di_erent patterns (1-4 shown below) to produce circles (used for the bin bases) as well as rectangles (that will form the bin sides). For example, if pattern 1 is used then 2 can sides and 2 circular bases are produced whilst pattern 3 produces 3 bin sides. Hynemann Co. has an order for 900 special edition cylindrical waste bins to be produced in one production run. Each bin requires one (1) circular base as well as one (1) rectangular side. Using a combination of the four patterns the company wants to produce the special edition bins in such a way that minimizes the number of raw metal sheets that are cut. Help Hynemann Co with this problem using linear programming. Provide a practical answer.

OpenStudy (anonymous):

the picture for the patterns is in the attachment . someone please help!!

OpenStudy (anonymous):

What I would do is create a problem with four variables Ax+By+Cz+Dw, where they are sheets one, two three, and four respectively

OpenStudy (anonymous):

I was thinking pattern one: 2x+2y<=900 Pattern 2: x+5y<=900 pattern 3: 3x<=900 pattern 4:6y<=900 and the objective function as minimise z=x+y Do you know much about linear programming ? I know its based around that

OpenStudy (anonymous):

I'm afraid I don't know of linear programming